Calculate the kinetic energy of a proton whose velocity is 0.8c.


1
Expert's answer
2021-04-12T06:55:59-0400

The relativistic kinetic energy of a proton is given as follows:


"E_k =\frac{m c^2}{\sqrt{1- v^2/c^2 }}-m c^2= mc^2\left( \frac{1}{\sqrt{1- v^2/c^2 }} - 1\right)"

where "mc^2 = 938\space MeV" is the rest energy of a proton, "v = 0.8c" is the speed of the proton, and "c" is the speed of light. Thus, obtain:


"E_k = 938\space MeV\left( \frac{1}{\sqrt{1- 0.8^2c^2/c^2 }} - 1\right) \approx 625.3\space MeV\approx 10^{-10}J"

Answer. "10^{-10}J".